# Canonical bases of tensor products of integrable highest weight modules arising from framed constructions

### v2

@@ 5.1. Kashiwara’s operators on $\mathbf{f}\theta_{\lambda}\mathbf{f}$
such that $b=\theta_{\lambda}b_{0}\in\theta_{j}\tilde{\mathbf{f}}$, and so $t_{j}(b)\geqslant 1$, a contradiction. Hence by [13, Section 14.3.1], we have $t_{i}^{\sigma}(\sigma(b_{0}))=t_{i}(b_{0})\leqslant\langle i,\lambda\rangle$ for any $i\in I$, and so $b_{0}\in\sigma(\mathbf{B}(\lambda))$. Therefore, the map $\theta_{\lambda}\cdot:\sigma(\mathbf{B}(\lambda))\rightarrow\bigcap_{i\in I}\tilde{\mathbf{B}}_{i,0}\cap\tilde{\mathbf{B}}(\mathbf{f}\theta_{\lambda}\mathbf{f})$ i surjective, and so it is a bijection. ∎
@@ 5.2. Kashiwara’s operators on $\Lambda_{\xi,\lambda}$
Let $L(\tilde{\Lambda}_{\xi\odot\lambda})=\{m\in\tilde{\Lambda}_{\xi\odot\lambda}\mid(m,m)_{\xi\odot\lambda}\in\mathbf{A}\}$ By Lemma 2.4 and Theorem 2.8, we know that $L(\tilde{\Lambda}_{\xi\odot\lambda})$ is the $\mathbf{A}$-submodule of $\tilde{\Lambda}_{\xi\odot\lambda}$ generated by $\tilde{\mathbf{B}}(\tilde{\Lambda}_{\xi\odot\lambda})$. We have the direct sum decomposition $L(\tilde{\Lambda}_{\xi\odot\lambda})=\bigoplus_{\tilde{\nu}\in\mathbb{N}[\tilde{I}]}L(\tilde{\Lambda}_{\xi\odot\lambda})_{\tilde{\nu}}$, where $L(\tilde{\Lambda}_{\xi\odot\lambda})_{\tilde{\nu}}$ is the $\mathbf{A}$-submodule generated by $\{\pi_{\xi\odot\lambda}(b)\mid b\in\tilde{\mathbf{B}}(\xi\odot\lambda)\cap\tilde{\mathbf{B}}_{\tilde{\nu}}\}$, see [13, Section 18.1.8].
@@
We know that $L^{\prime}(\Lambda_{\xi,\lambda})$ is a submodule of $L(\Lambda_{\xi,\lambda})$. For any $\nu\in\mathbb{N}[I]$, we make an induction on $\mathrm{tr}\,|\nu|\in\mathbb{N}$ to prove that
@@
We make an induction on $\mathrm{tr}\,\nu$ to prove the statement.
@@
$=$ $(\varphi(L^{\prime}(\Lambda_{\xi,\lambda})^{\nu-i-|b_{0}|,b_{0}}),\varphi\tilde{E}_{i}(L^{\prime}(\Lambda_{\xi,\lambda})^{\nu-|b^{\prime}_{0}|,b^{\prime}_{0}}))_{\xi,\lambda}\ \mathrm{mod}\,v^{-1}\mathbf{A}$

### v3

@@ 5.1. Kashiwara’s operators on $\mathbf{f}\theta_{\lambda}\mathbf{f}$
such that $b=\theta_{\lambda}b_{0}\in\theta_{j}\tilde{\mathbf{f}}$, and so $t_{j}(b)\geqslant 1$, a contradiction. Hence by [13, Section 14.3.1], we have $t_{i}^{\sigma}(\sigma(b_{0}))=t_{i}(b_{0})\leqslant\langle i,\lambda\rangle$ for any $i\in I$, and so $b_{0}\in\sigma(\mathbf{B}(\lambda))$. Therefore, the map $\theta_{\lambda}\cdot:\sigma(\mathbf{B}(\lambda))\rightarrow\bigcap_{i\in I}\tilde{\mathbf{B}}_{i,0}\cap\tilde{\mathbf{B}}(\mathbf{f}\theta_{\lambda}\mathbf{f})$ is surjective, and so it is a bijection. ∎
Similarly, there is a bijection
(14) $\cdot\theta_{\lambda}:\mathbf{B}(\lambda)\rightarrow\bigcap_{i\in I}\tilde{\mathbf{B}}_{i,0}^{\sigma}\cap\tilde{\mathbf{B}}(\mathbf{f}\theta_{\lambda}\mathbf{f}).$
@@ 5.2. Kashiwara’s operators on $\Lambda_{\xi,\lambda}$
Let $L(\tilde{\Lambda}_{\xi\odot\lambda})=\{m\in\tilde{\Lambda}_{\xi\odot\lambda}\mid(m,m)_{\xi\odot\lambda}\in\mathbf{A}\}$ . By Lemma 2.4 and Theorem 2.8, we know that $L(\tilde{\Lambda}_{\xi\odot\lambda})$ is the $\mathbf{A}$-submodule of $\tilde{\Lambda}_{\xi\odot\lambda}$ generated by $\tilde{\mathbf{B}}(\tilde{\Lambda}_{\xi\odot\lambda})$. We have the direct sum decomposition $L(\tilde{\Lambda}_{\xi\odot\lambda})=\bigoplus_{\tilde{\nu}\in\mathbb{N}[\tilde{I}]}L(\tilde{\Lambda}_{\xi\odot\lambda})_{\tilde{\nu}}$, where $L(\tilde{\Lambda}_{\xi\odot\lambda})_{\tilde{\nu}}$ is the $\mathbf{A}$-submodule generated by $\{\pi_{\xi\odot\lambda}(b)\mid b\in\tilde{\mathbf{B}}(\xi\odot\lambda)\cap\tilde{\mathbf{B}}_{\tilde{\nu}}\}$, see [13, Section 18.1.8].
@@
We know that $L^{\prime}(\Lambda_{\xi,\lambda})$ is a submodule of $L(\Lambda_{\xi,\lambda})$. For any $\nu\in\mathbb{N}[I]$, we use induction on $\mathrm{tr}\,|\nu|\in\mathbb{N}$ to prove that
@@
We use induction on $\mathrm{tr}\,\nu$ to prove the statement.
@@
$=$ $(\varphi(L^{\prime}(\Lambda_{\xi,\lambda})^{\nu-i-|b_{0}|,b_{0}}),\varphi\tilde{E}_{i}(L^{\prime}(\Lambda_{\xi,\lambda})^{\nu-|b^{\prime}_{0}|,b^{\prime}_{0}}))_{\xi,\lambda}\ \mathrm{mod}\,v^{-1}\mathbf{A}.$